OpenStudy (anonymous):

Simplify: 4sqrt400/4sqrt5 Please do not just give out the answer I need someone to show work so I can understand

OpenStudy (mrnood):

ok first of all do you understand that \[\sqrt{ab} = \sqrt{a} \times \sqrt{b}\] ?

OpenStudy (anonymous):

Yea I think

OpenStudy (johnweldon1993):

\[\frac{4\sqrt{400}}{4\sqrt{5}}\] First we can cancel the 4's...since we have a 4 on both top and bottom they will divide to 1 and cancel...so \[\large \frac{\sqrt{400}}{\sqrt{5}}\] Now we know \[\large \frac{\sqrt{a}}{\sqrt{b}} = \sqrt{\frac{a}{b}}\] so now we can write this as \[\large \sqrt{\frac{400}{5}}\] Just do the division and simplify a bit :)

OpenStudy (anonymous):

Thanks @johnweldon1993 that actually helped I got how to do it now